# A Bayesian Flow Network Framework for Chemistry Tasks

### v1

@@ 3.2 Conditional Generation of Small Molecules
The classifier-free guidance32 method is easily adapted into BFN, where only the computing of output distribution needs changing during sampling process. The pseudocode of computing discrete output distribution is presented in Algorithm 1. In the experiment, we jointly trained a model conditional and unconditional on QM9 dataset with an unconditional rate $p_{uncond}=0.2$. In the sampling stage, $w$ was set to 4. We sampled 10 molecules using the label [-0.249, 0.0615, 0.3105] that was transformed to $\boldsymbol{y}$ via a trained 2-layer MLP. 10 unconditioned samples were generated as a control group. RDKit33 was employed to generate the 3D conformations, then the geometry optimisations and energy calculations were performed via PySCF34 at B3LYP/6-31G(2df,p) level of accuracy. The results of MAE between calculated values and labels are presented in Table 6. The conditioned samples are displayed in Figure 5.
@@
Table 6: MAE on QM9 dataset w/ and w/o classifier-free guidance generation. Smaller errors are in bold.
@@ 3.3 Molecular Scaffold Extension
Here, we show a simple inpaint strategy can extend molecular scaffolds by using ChemBFN. In every sampling steps, parameters of input distributions are modified as $\boldsymbol{\theta}\leftarrow\boldsymbol{M}\odot\boldsymbol{e_{x}}+(1-\boldsymbol{M})\odot\boldsymbol{\theta}$ before being inputted into the network, where $\boldsymbol{M}$ is the mask and $\boldsymbol{e_{x}}$ is the one-hot representation of scaffold. Figure 6 shows an example of extending scaffold ‘Cc1cc(OC5)cc(C6)c1.’ by a model trained on MOSES SAFE35 version, a variation of SMILES. We found that inpainting sampling for 10 to 100 steps was sufficient to generate complex molecules.
@@ 3.4 Finetuning on Prediction Tasks
In this section, we compare our model with SOTA models36, 37, 38, 39, 40, 30, 31, 41, including graph-based and language-based which could be further classified as smaller scale natural language processing models (NLPs) and large language models (LLMs), on subsets of MoleculeNet benchmark. As shown in Table 7, our method outperformed SOTA language models on several tasks, especially ClinTox and BBBP. It is notable that ChemBERTa30 and ChemBERTa-231, which had a similar model size with ours, were pretrained on 77M molecules but had worse scores on 3 out of 5 tasks than ours. This indicated that BFN-style generative pretraining is a better strategy than masked language modeling and multitask regression pretraining. A similar observation applied to CaRRoBERTa model that coupled the knowledge of ChatGPT42 (which is far larger in scale than ours and is believed to have seen more chemical texts) and the distillation capability of RoBERTa21 method: our model outperformed CaRRoBERTa on 4 out of 5 tasks. However, when comparing with graph neural networks (GNNs) our model performed averagely 1.7% worse, especially on regression tasks.
Table 7: Testing metrics on sub-tasks of MoleculeNet benchmark with scaffold splitting compared with SOTA models. The metrics of all other models were copied from their original paper. $\uparrow$ indicates that the higher is better and $\downarrow$ stands for the contrary. The best results are in bold. The best results within the same category (graph-based or language-based) are underlined. Percentages in the last two rows show the performance changes w.r.t the best models and the colour represents whether our model was better (in red) or not (in blue).
Method ROC-AUC $\uparrow$ RMSE $\downarrow$
ClinTox BBBP BACE HIV ESOL FreeSolv Lipo
GNNs Uni-Mol36 91.9 72.9 85.7 80.8 0.788 1.480 0.603
MolKD37 83.8 74.8 80.1 74.9
GEM38 90.1 72.4 85.6 80.6 0.798 1.877 0.660
Mole-BERT39 78.9 71.9 80.8 78.2 1.015 0.676
LLMs CaRRoBERTa40 84.16 81.99 80.73 0.96 1.02
NLPs ChemBERTa30 73.3 64.3 62.2
ChemBERTa-231 60.1 74.2 79.9 0.744
SMILES Transformer41 70.4 70.1 72.9
ChemBFN (ours) 99.18 95.74 73.56 79.37 0.884 1.418 0.746
$\Delta_{GNNs_{best}}$ +8% +28% -14% -2% +12% -4% +24%
$\Delta_{LMs_{best}}$ +18% +17% -9% +9% -8% 0%
@@ 3.5 Reaction Yield Prediction
In order to predict the reaction yield, we first trained the generative model to understand chemical reaction by learning to predict the products. We developed an in-context style guidance that during training stage only the parameters of product in reaction SMILES were predicted by always masking the input distribution of reactant/reagent and >>much-greater-than$>$ > token to the one-hot representation, i.e., $\boldsymbol{\theta}\leftarrow\boldsymbol{M}_{rr}\odot\boldsymbol{e_{x}}+(1-\boldsymbol{M}_{rr})\odot\boldsymbol{\theta}$, where $\boldsymbol{M}_{rr}$ is the mask for reactant, reagent, and >>much-greater-than$>$ > token.
The generative model was first pre-trained on USPTO-50k dataset then post-trained on Buchwald-Hartwig and Suzuki-Miyaura coupling datasets before the whole prediction model was fine-tuned. The testing scores compared with previous researches43, 28, 44 were reported in Table 8. It is notable that the Yield-BERT series28, 44 were based on a pre-trained RXNFP45 model which had been pre-trained on over 2M reactions while our model was pre-trained on 50k reactions. Despite the disadvantage of limited access of pretraining data, the performance of our method was still close to that of largely pretrained model on random-split sets and significantly better on out-of-sample predictions.
Table 8: R2 scores on different testing sets of HTE Buchwald-Hartwig and Suzuki-Miyaura reaction datasets. The scores of all other models were copied from original paper. The best results are in bold. The score of “rand 70/30” split was the 10-fold average value. Test 1-4 were out-of-sample splits.
@@
Dataset Split MFF43 Yield-BERT28 Yield-BERT-DA44 ChemBFN (ours)
@@
Test 1 0.85 0.84 0.82 0.844
Test 2 0.71 0.84 0.90 0.910
Test 3 0.64 0.75 0.63 0.787
Test 4 0.18 0.49 0.43 0.633
Avg. 1-4 0.60 0.73 0.69 0.794
@@ 3.6 Is Larger Pretrain Dataset Better?
We have seen that our model, although was pretrained on 40M molecules, outperformed models pretrained on larger dataset on several prediction tasks. Here rises a question: whether larger pretrain dataset benefits our method? To answer this, 3 models were trained on AqSolDB dataset, of which one was trained from scratch, one was pretrained on 40M molecules from ZINC15 database, and one was pretrained on 190M molecules from ZINC15. We summarised the testing results in Table 9. Interestingly, the errors did not shrink when the pretrain data grew from 40M to 190M. However, compared with zero pretraining, an improvement on performance of $\geq$12.5can be confirmed.
Table 9: Testing metrics of models with different pretrain data sizes (0, 40M, and 190M) on AqSolDB dataset.
@@ 3.7 Training Details
During fine-tuning stages, models were trained for 100 epochs on labelled datasets. The batch-size, both for training and validation, was 32 on MoleculeNet benchmark; the training batch-size was 16 for reaction yield prediction. $lr_{max}$ was $10^{-4}$ that was warmed up from $10^{-7}$ during the first 1,000 steps for regression tasks and 100 steps for classification tasks. After warm-up stage, $lr$ decreased by 0.2 after the validation metrics stopped improving for 20 epochs unless the learning rate had reached $10^{-6}$. The dropout rate of prediction MLP head was fine-tuned for each case and we recommend to try from $\{0.0,0.5,0.7\}$. The validation metrics for regression and classification tasks were MAE and accuracy, respectively.
We employed AdamW46 with default hyperparameters implemented in PyTorch47 as the optimizer for all tasks.
@@ 5 Acknowledgements
We show gratefulness to Research Center for Computational Science (RCCS), Okazaki, Japan and its maintenance team, who provided computing sources including A100 GPUs. This work is under project 24-IMS-C043 of RCCS. We also thank Dr. Maho Nakata who kindly lent us his own RTX 3080 GPU. Without the kindness of above mentioned people and team, this work would never be finished.

### v2

@@ 3.2 Conditional Generation of Small Molecules
The classifier-free guidance45 method is easily adapted into BFN, where only the computation of output distribution needs changing during sampling process. The pseudocode for computing discrete output distribution is presented in Algorithm 1. In the experiment, we jointly trained a model conditionally and unconditionally on QM9 dataset with an unconditional rate $p_{uncond}=0.2$. In the sampling stage, $w$ was set to 4. We sampled 10 molecules using the label [-0.249, 0.0615, 0.3105] that was transformed to $\boldsymbol{y}$ via a trained 2-layer MLP. 10 unconditioned samples were generated as a control group. RDKit46 was employed to generate the 3D conformations, then the geometry optimisations and energy calculations were performed via PySCF47 at B3LYP/6-31G(2df,p) level of accuracy. The results of MAE between calculated values and labels are presented in Table 8. The conditioned samples are displayed in Figure 5.
@@
Table 8: MAE on QM9 dataset w/ and w/o classifier-free guidance generationa
@@ 3.3 Molecular Scaffold Extension
Here, we show a simple inpaint strategy can extend molecular scaffolds by using ChemBFN. In every sampling steps, parameters of input distributions are modified as $\boldsymbol{\theta}\leftarrow\boldsymbol{M}\odot\boldsymbol{e_{x}}+(1-\boldsymbol{M})\odot\boldsymbol{\theta}$ before being inputted into the network, where $\boldsymbol{M}$ is the mask and $\boldsymbol{e_{x}}$ is the one-hot representation of scaffold. Figure 6 shows an example of extending scaffold ‘Cc1cc(OC5)cc(C6)c1.’ by a model trained on MOSES SAFE48 version, a variation of SMILES. We found that inpainting sampling for 10 to 100 steps was sufficient to generate complex molecules.
@@ 3.4 Finetuning on Prediction Tasks
In this section, we compare our model with SOTA models49, 50, 51, 52, 53, 43, 44, 54, 55, including graph-based and language-based which could be further classified as smaller scale natural language processing models (NLPs) and large language models (LLMs), on subsets of MoleculeNet benchmark. As shown in Table 9, our method outperformed SOTA language models on several tasks, especially ClinTox and BBBP. It is notable that ChemBERTa43 and ChemBERTa-244, which had a similar model size with ours, were pretrained on 77M molecules but had worse scores on 3 out of 5 tasks than ours. This indicated that BFN-style generative pretraining is a better strategy than masked language modeling and multitask regression pretraining. A similar observation applied to CaRRoBERTa model that coupled the knowledge of ChatGPT56 (which is far larger in scale than ours and is believed to have seen more chemical texts) and the distillation capability of RoBERTa29 method: our model outperformed CaRRoBERTa on 4 out of 5 tasks. However, when comparing with graph neural networks (GNNs) our model performed averagely 1.7% worse, especially on regression tasks.
Table 9: Testing metrics on sub-tasks of MoleculeNet benchmark with scaffold splitting compared with SOTA modelsa
[…]
[][c] Method ROC-AUC $\uparrow$ RMSE $\downarrow$ ClinTox BBBP BACE HIV ESOL FreeSolv Lipo GNNs Uni-Mol49 91.9 $\pm$ 1.8 72.9 $\pm$ 0.6 85.7 $\pm$ 0.2 80.8 $\pm$ 0.3 0.788 $\pm$ 0.029 1.480 $\pm$ 0.048 0.603 $\pm$ 0.010 MolKD50 83.8 $\pm$ 3.1 74.8 $\pm$ 2.3 80.1 $\pm$ 0.8 74.9 $\pm$ 1.7 GEM51 90.1 $\pm$ 1.3 72.4 $\pm$ 0.4 85.6 $\pm$ 1.1 80.6 $\pm$ 0.9 0.798 $\pm$ 0.029 1.877 $\pm$ 0.094 0.660 $\pm$ 0.008 Mole-BERT52 78.9 $\pm$ 3.0 71.9 $\pm$ 1.6 80.8 $\pm$ 1.4 78.2 $\pm$ 0.8 1.015 $\pm$ 0.030 0.676 $\pm$ 0.017 LLMs CaRRoBERTa53 84.16 $\pm$ 17.63 81.99 $\pm$ 4.19 80.73 $\pm$ 1.42 0.96 $\pm$ 0.09 1.02 $\pm$ 0.06 NLPs ChemBERTa43 73.3 64.3 62.2 ChemBERTa-244 60.1 74.2 79.9 0.744 AGBT55 76.3 SMILES Transformer54 70.4 70.1 72.9 ChemBFN (ours) 99.18 $\pm$ 1.77 95.74 $\pm$ 0.70 73.56 $\pm$ 1.22 79.37 $\pm$ 1.66 0.884 $\pm$ 0.003 1.418 $\pm$ 0.067 0.746 $\pm$ 0.001 $\Delta_{GNNs_{best}}$ +8% +28% -14% -2% +12% -4% +24% $\Delta_{LMs_{best}}$ +18% +17% -9% +9% -8% 0%
  • a
    The metrics of all other models were copied from their original paper. $\uparrow$ indicates that the higher is better and $\downarrow$ stands for the contrary. The best results are in bold. The best results within the same category (graph-based or language-based) are underlined. Percentages in the last two rows show the performance changes w.r.t the best models and the colour represents whether our model was better (in red) or not (in blue).
We further benchmarked our model on the public ADME dataset38 (regression task) and Kinase inhibitor dataset39 (classification task). The results for the public ADME dataset were summarised in Table 10. For the Kinase inhibitor dataset, the averaged ROC-AUC over 354 assays tested on the random split was (87.93 $\pm$ 14.05)and the averaged ROC-AUC tested on the scaffold split was (79.35 $\pm$ 18.71)%.
Table 10: MAE, RMSE, and Pearson’s correlation coefficient on the public ADME dataset.
HLM RLM hPPB rPPB MDR1-MDCK ER Solubility
MAE 0.359 $\pm$ 0.005 0.428 $\pm$ 0.001 0.365 $\pm$ 0.006 0.408 $\pm$ 0.006 0.337 $\pm$ 0.003 0.411 $\pm$ 0.008
RMSE 0.474 $\pm$ 0.004 0.556 $\pm$ 0.005 0.479 $\pm$ 0.012 0.549 $\pm$ 0.009 0.466 $\pm$ 0.007 0.630 $\pm$ 0.013
R 0.653 $\pm$ 0.002 0.685 $\pm$ 0.007 0.771 $\pm$ 0.001 0.700 $\pm$ 0.003 0.765 $\pm$ 0.007 0.588 $\pm$ 0.003
@@ 3.5 Reaction Yield Prediction
In order to predict the reaction yield, we first trained the generative model to understand chemical reaction by learning to predict the products. We developed an in-context style guidance that during training stage, only the parameters of product in reaction SMILES were predicted. This was achieved by always masking the input distribution of reactant/reagent and >>much-greater-than$>$ > tokens that were converted to the corresponding one-hot representation, i.e., $\boldsymbol{\theta}\leftarrow\boldsymbol{M}_{rr}\odot\boldsymbol{e_{x}}+(1-\boldsymbol{M}_{rr})\odot\boldsymbol{\theta}$, where $\boldsymbol{M}_{rr}$ is the mask for reactant, reagent, and >>much-greater-than$>$ > token.
The generative model was first pre-trained on USPTO-50k dataset then post-trained on Buchwald-Hartwig and Suzuki-Miyaura coupling datasets before the whole prediction model was fine-tuned. The testing scores compared with previous researches57, 41, 58 were reported in Table 11. It is notable that the Yield-BERT series41, 58 were based on a pre-trained RXNFP59 model which had been pre-trained on over 2M reactions while our model was pre-trained on 50k reactions. Despite the disadvantage of limited access of pretraining data, the performance of our method was still close to that of largely pretrained model on random-split sets and significantly better on out-of-sample predictions.
Table 11: R2 scores on different testing sets of HTE Buchwald-Hartwig and Suzuki-Miyaura reaction datasetsa
@@
Dataset Split MFF57 Yield-BERT41 Yield-BERT-DA58 ChemBFN (ours)
@@
Test 1 0.85 0.84 $\pm$ 0.010 0.82 $\pm$ 0.01 0.844 $\pm$ 0.002
Test 2 0.71 0.84 $\pm$ 0.03 0.90 $\pm$ 0.01 0.910 $\pm$ 0.001
Test 3 0.64 0.75 $\pm$ 0.04 0.63 $\pm$ 0.05 0.787 $\pm$ 0.034
Test 4 0.18 0.49 $\pm$ 0.05 0.43 $\pm$ 0.07 0.633 $\pm$ 0.082
Avg. 1-4 0.60 0.73 $\pm$ 0.15 0.69 $\pm$ 0.19 0.794 $\pm$ 0.118
@@
  • a
    The scores of all other models were copied from the original paper. The best results are in bold. The score of “rand 70/30” split was the 10-fold average value. Test 1-4 were out-of-sample splits.
@@ 3.6 Is Larger Pretrain Dataset Better?
We have seen that our model, although was pretrained on 40M molecules, outperformed the models pretrained on larger dataset on several prediction tasks. Here rises a question: does a larger pretraining dataset benefit our method? To answer this, three models were trained on AqSolDB dataset, of which one was trained from scratch, one was pretrained on 40M molecules from ZINC15 database, and the third one was pretrained on 190M molecules from ZINC15. We summarised the testing results in Table 12. Interestingly, the errors did not shrink when the pretraining data grew from 40M to 190M. However, compared with zero pretraining, an improvement in performance of $\geq$12.5can be confirmed.
Table 12: Testing metrics of models with different pretrain data sizes (0, 40M, and 190M) on AqSolDB dataset
@@ 3.7 Training Details
During fine-tuning stages, models were trained for 100 epochs on labelled datasets. The batch-size, both for training and validation, was 32 on MoleculeNet benchmark, AqSolDB dataset, public ADME dataset, and Kinase inhibitor dataset; the training batch-size was 16 for reaction yield prediction. $lr_{max}$ was $10^{-4}$ that was warmed up from $10^{-7}$ during the first 1,000 steps for regression tasks and 100 steps for classification tasks. After warm-up stage, $lr$ decreased by 0.2 after the validation metrics stopped improving for 20 epochs unless the learning rate had reached $10^{-6}$. The dropout rate of prediction MLP head was fine-tuned for each case and we recommend to try from $\{0.0,0.1,0.5,0.7\}$. The validation metrics for regression and classification tasks were MAE and inverted accuracy (i.e., 1 - accuracy), respectively.
We employed AdamW60 with default hyperparameters implemented in PyTorch61 as the optimizer for all tasks.
@@ 4 Conclusion
5 Data and Software Availability
The code, pre-trained models, and instructions necessary to reproduce the results of this study are available for download at https://github.com/Augus1999/bayesian-flow-network-for-chemistry.
@@ 6 Acknowledgements
We express our gratitude to the Research Center for Computational Science (RCCS) in Okazaki, Japan, and its maintenance team for providing computing resources, including A100 GPUs. This work is under project 24-IMS-C043 of RCCS. We also thank Dr. Maho Nakata who kindly lent us his own RTX 3080 GPU and Prof. Kazumasa Okada for discussion.
7 Conflict of Interest
There is no conflict of interest.
8 Funding Sources
The authors claim that there is no funding related to this research.